6. XYZ company’s cost function for the next four months is C = 500,000+5Q


(a) Find the BE dollar volume of sales if the selling price is br. 6 / unit


(b) What would be the company’s cost if it decides to shut down operations for the next four months


(c) If, because of strike, the most the company can produce is br. 100,000 units, should it


shutdown? Why or why not?

1
Expert's answer
2022-04-11T11:30:37-0400

The selling price is given as 6/unit.

A.

If the number of units are 'Q' then the total selling price is equal to (6×Q) 

(Volume of sales) = (Selling price)-(Cost price) 

"=(6×Q) -(500000+5Q) \\

 =6Q-500000-5Q\\

 =Q-500000\\"


B.

If the company decides to shut down for next four months, we know that there will be no units left. That means Q=0

"C=500000+5Q\\

C=500000+5(0) \\

C=500000"

Therefore the cost of the company will be 500000


C.


From the task 1 we know that the volume of sales"=(Q-500000)"  

If the company produces 100000 units then the sales will be "(100000-500000) =-400000"

Negative indicates the $400000 loss. 

Hence the company should shut down as it in loss.